helpmenow88:

One pipe fills a storage pool in 16 hours. A second pipe fills the same pool in 24 hours. When a third pipe is added and all three are used to fill the​ pool, it takes only 6 hours. Find how long it takes the third pipe to do the job alone.

helpmenow88:

help

justjm:

This type of question is a classic in the algebra 2 level. Think about it this way; if one pipe fills the storage pool in 16 hours, that means that in 1 hour, 1/16 of the job is done. Similarly, for the second pipe, it would be 1/24 of the job, and for the third pipe, it is 1/xth of the job. When they are all put to the job, in one hour, 1/6th of the job is completed. Now you can form an equation resembling this scenario \(\frac{1}{16} + \frac{1}{24} + \frac{1}{x}=\frac{1}{6} \)

justjm:

And then you can solve for x

helpmenow88:

thanks!

justjm:

Np, let me know if you have any questions